Find two positive integers whose difference is 4 and whose product is 96
Answer by checkley77(12844)

Let x & x=4 be the 2 integers.
x(x+4)=96
x^2+4x-96=0
(x-8)(x=12)=0
x-8=0
x=8 ans. for the smaller integer.
8+4=12for the larger integer.
Proof:
8*12=96
96=96